# Video Repurposing from User Generated Content: A Large-scale Dataset and Benchmark

### v1

@@ Abstract
The demand for producing short-form videos for sharing on social media platforms has experienced significant growth in recent times. Despite notable advancements in the fields of video summarization and highlight detection, which can create partially usable short films from raw videos, these approaches are often domain-specific and require an in-depth understanding of real-world video content. To tackle this predicament, we propose Repurpose-10K, an extensive dataset comprising over 10,000 videos with more than 120,000 annotated clips aimed at resolving the video long-to-short task. Recognizing the inherent constraints posed by untrained human annotators, which can result in inaccurate annotations for repurposed videos, we propose a two-stage solution to obtain annotations from real-world user-generated content. Furthermore, we offer a baseline model to address this challenging task by integrating audio, visual, and caption aspects through a cross-modal fusion and alignment framework. We aspire for our work to ignite groundbreaking research in the lesser-explored realms of video repurposing. The code and data will be available at https://github.com/yongliang-wu/Repurpose.
@@ Introduction
Driven by the rapid growth of social media platforms such as Instagram, TikTok, and YouTube Shorts, short-form videos become the primary medium for sharing daily life and conveying information (Xiong et al. 2019). Even creators who traditionally focus on long-form content, such as live streams, interviews, and vlogs, now shift towards producing captivating short-form videos for these platforms (Wang et al. 2023b). This shift underscores the importance of identifying the most engaging clips within longer videos while maintaining logical coherence.
@@
Temporal Event Localization (TEL) (Tian et al. 2018; Tian, Li, and Xu 2020; Lee et al. 2021; Geng et al. 2023; Zhang et al. 2024; Li et al. 2024a; Zhang, Zhang, and Zhou 2024) is a video understanding task that aims to identify and pinpoint the start and end times of events or actions of interest within untrimmed videos. This task proves crucial for areas such as multimedia content analysis, video surveillance, and sports event analysis. The supervised learning paradigms of TEL are categorized into two primary methodologies: two-stage approaches (Shou, Wang, and Chang 2016; Zhao et al. 2017; Lin et al. 2018; Li et al. 2023a, 2024c) and single-stage frameworks (Buch et al. 2019; Zhang, Wu, and Li 2022; Yan et al. 2023; Li et al. 2024b; Liu, Li, and Yu 2024). Recently, UnLoc (Yan et al. 2023) proposes a unified framework for video localization tasks, consisting of a two-tower CLIP model, with output features fed into a video-text fusion module and feature pyramid. Unlike TEL, which typically focuses on identifying explicit timestamps, video repurposing revolves around identifying complete and engaging segments within lengthy videos, without explicitly marking a specific starting point, as a temporal localization task (Liu et al. 2024).
@@
Video highlight detection (Sun, Farhadi, and Seitz 2014; Xiong et al. 2019) aims to automatically identify the most compelling and quintessential segments of a video. While earlier methods consider only visual features (Sun, Farhadi, and Seitz 2014), recent works start to incorporate both audio and visual components to retrieve highlights from the video (Badamdorj et al. 2021). Query-Dependent DETR (Moon et al. 2023) employs a cross-attention transformer encoder to generate more potent multi-modal video representations, deliberately infusing text queries into the feature representation. However, highlight detection particularly emphasizes moments that are exceptionally striking or exhilarating, rather than necessarily providing a coherent summary of the video content. For example, a moment is considered a highlight if at least 70% of its frames match the edited video in the training of YouTube Highlights (Sun, Farhadi, and Seitz 2014). This approach to combining highlights poses challenges in crafting a comprehensive narrative thread, making it difficult for direct publication on short-form social media platforms due to the lack of a cohesive storyline. Additionally, repurposing video highlights often requires post-processing to form coherent segments, introducing further effort.
@@
Video summarization (De Avila et al. 2011; Gygli et al. 2014; Song et al. 2015) aims to extract key information from lengthy video content and generate a concise summary version. This task preserves the most informative and representative parts of the video, allowing users to quickly grasp the main content without watching the entirety of the original video. The key parts within a video could either be a series of key frames or a collection of one or more key segments. To better capture temporal correlations in video content, a variety of sophisticated deep learning models (Zhang et al. 2016; Li et al. 2022) develop to map dependencies across time, using either localized or holistic approaches. Video summarization prioritizes comprehensiveness, often overlooking the engaging aspect of video content. In contrast, video repurposing does not mandate conveying the entire essence of the video.
@@
We approach video repurposing by jointly learning classification and regression tasks, taking video elements (segments, audio, captions) as input and outputting the time range for selected clips. The classification task determines if a segment should be selected, while the regression task pinpoints the exact temporal offsets for the clip’s start and end. For an input video sequence with visual and audio tracks, we segment it into visual and audio pairs, represented as $\{V_{t},A_{t}\}_{t=1}^{T}$, where $T$ is the number of segments. In cases involving human speech, we use whisperX (Bain et al. 2023) to convert audio segments into time-stamped captions, $C_{t}$, enhancing our segment pairs to $\{V_{t},A_{t},C_{t}\}_{t=1}^{T}$. For each video, we select a set of clips, compiling temporal intervals denoted by $R=\{(t_{s,n},t_{e,n})\}_{n=1}^{N}$, where $N$ is the number of curated excerpts.
@@
To further demonstrate the effectiveness of our model, we present two practical applications as illustrated in Figure 4. In both instances, there is a notable alignment between our model’s predictions and the actual ground truth. In the first scenario, the ground truth annotation provides a straightforward and concise description, leading to the decision to explore these “strange places.” In contrast, our model’s prediction starts with an engaging B-roll news clip and ends with the main character’s declaration, “Let the adventure begin.” In the second scenario, while the annotation offers a detailed introduction to the product of the video creator, our predictions more effectively highlight the video’s concluding section. Both selected clips begin with intriguing content and end with fitting remarks, demonstrating their ability to engage and provide a complete narrative. This result indicates that our model achieves a comprehensive understanding of video content, enabling it to select clips suitable for the video repurposing task.
@@ Conclusion
We curate the Repurpose-10K dataset specifically for the challenging task of video repurposing. The dataset comprises over 10,000 videos and more than 120,000 clips, annotated with User Generated Content (UGC) in the real world. To validate the utility of UGC annotations, we conduct a comprehensive double-blind user study comparing video repurposing outcomes using UGC annotations with those of randomly selected clips and clips edited by professional video editors. Furthermore, we present an end-to-end model designed to address the video repurposing challenge by jointly learning classification and regression tasks. Through extensive experiments and ablation studies, we demonstrate the effectiveness of the proposed model.

### v2

@@ Abstract
The demand for producing short-form videos for sharing on social media platforms has experienced significant growth in recent times. Despite notable advancements in the fields of video summarization and highlight detection, which can create partially usable short films from raw videos, these approaches are often domain-specific and require an in-depth understanding of real-world video content. To tackle this predicament, we propose Repurpose-10K, an extensive dataset comprising over 10,000 videos with more than 120,000 annotated clips aimed at resolving the video long-to-short task. Recognizing the inherent constraints posed by untrained human annotators, which can result in inaccurate annotations for repurposed videos, we propose a two-stage solution to obtain annotations from real-world user-generated content. Furthermore, we offer a baseline model to address this challenging task by integrating audio, visual, and caption aspects through a cross-modal fusion and alignment framework. We aspire for our work to ignite groundbreaking research in the lesser-explored realms of video repurposing.
Codehttps://github.com/yongliang-wu/Repurpose
@@ Introduction
Driven by the rapid growth of social media platforms such as Instagram, TikTok, and YouTube Shorts, short-form videos become the primary medium for sharing daily life and conveying information. Even creators who traditionally focus on long-form content, such as live streams, interviews, and vlogs, now shift towards producing captivating short-form videos for these platforms (Wang et al. 2023b). This shift underscores the importance of identifying the most engaging clips within longer videos while maintaining logical coherence.
@@
Temporal Event Localization (TEL) (Tian et al. 2018; Tian, Li, and Xu 2020; Lee et al. 2021; Geng et al. 2023; Zhang et al. 2024; Li et al. 2024a; Zhang, Zhang, and Zhou 2024) is a video understanding task that aims to identify and pinpoint the start and end times of events or actions of interest within untrimmed videos. This task proves crucial for areas such as multimedia content analysis, video surveillance, and sports event analysis. The supervised learning paradigms of TEL are categorized into two primary methodologies: two-stage approaches (Zhao et al. 2017; Li et al. 2023a, 2024c) and single-stage frameworks (Buch et al. 2019; Zhang, Wu, and Li 2022; Yan et al. 2023; Li et al. 2024b; Liu, Li, and Yu 2024). Recently, UnLoc (Yan et al. 2023) proposes a unified framework for video localization tasks, consisting of a two-tower CLIP model, with output features fed into a video-text fusion module and feature pyramid. Unlike TEL, which typically focuses on identifying explicit timestamps, video repurposing revolves around identifying complete and engaging segments within lengthy videos, without explicitly marking a specific starting point, as a temporal localization task (Liu et al. 2024).
@@
Video highlight detection (Sun, Farhadi, and Seitz 2014) aims to automatically identify the most compelling and quintessential segments of a video. While earlier methods consider only visual features (Sun, Farhadi, and Seitz 2014), recent works start to incorporate both audio and visual components to retrieve highlights from the video (Badamdorj et al. 2021). Query-Dependent DETR (Moon et al. 2023) employs a cross-attention transformer encoder to generate more potent multi-modal video representations, deliberately infusing text queries into the feature representation. However, highlight detection particularly emphasizes moments that are exceptionally striking or exhilarating, rather than necessarily providing a coherent summary of the video content. For example, a moment is considered a highlight if at least 70% of its frames match the edited video in the training of YouTube Highlights (Sun, Farhadi, and Seitz 2014). This approach to combining highlights poses challenges in crafting a comprehensive narrative thread, making it difficult for direct publication on short-form social media platforms due to the lack of a cohesive storyline. Additionally, repurposing video highlights often requires post-processing to form coherent segments, introducing further effort.
@@
Video summarization (Song et al. 2015) aims to extract key information from lengthy video content and generate a concise summary version. This task preserves the most informative and representative parts of the video, allowing users to quickly grasp the main content without watching the entirety of the original video. The key parts within a video could either be a series of key frames or a collection of one or more key segments. To better capture temporal correlations in video content, a variety of sophisticated deep learning models (Zhang et al. 2016; Li et al. 2022) develop to map dependencies across time, using either localized or holistic approaches. Video summarization prioritizes comprehensiveness, often overlooking the engaging aspect of video content. In contrast, video repurposing does not mandate conveying the entire essence of the video.
@@
We approach video repurposing by jointly learning classification and regression tasks, taking video elements (segments, audio, captions) as input and output the time range for selected clips. The classification task determines if a segment should be selected, while the regression task pinpoints the exact temporal offsets for the clip’s start and end. For an input video sequence with visual and audio tracks, we segment it into visual and audio pairs, represented as $\{V_{t},A_{t}\}_{t=1}^{T}$, where $T$ is the number of segments. In cases involving human speech, we use whisperX (Bain et al. 2023) to convert audio segments into time-stamped captions, $C_{t}$, enhancing our segment pairs to $\{V_{t},A_{t},C_{t}\}_{t=1}^{T}$. For each video, we select a set of clips, compiling temporal intervals denoted by $R=\{(t_{s,n},t_{e,n})\}_{n=1}^{N}$, where $N$ is the number of curated excerpts.
@@
To further demonstrate the effectiveness of our model, we present two practical applications as illustrated in Figure 4. In both instances, there is a notable alignment between our model’s predictions and the actual ground truth. In the first scenario, the ground truth annotation provides a straightforward and concise description, leading to the decision to explore these “strange places.” In contrast, our model’s prediction starts with an engaging B-roll news clip and ends with the main character’s declaration, “Let the adventure begin.” In the second scenario, while the annotation offers a detailed introduction to the product of the video creator, our predictions more effectively highlight the video’s concluding section. Both selected clips begin with intriguing content and end with fitting remarks, demonstrating their ability to engage and provide a complete narrative.
@@ Conclusion
In this paper, we curate the Repurpose-10K dataset for the challenging task of video repurposing. The dataset comprises over 10,000 videos and more than 120,000 clips, annotated with User Generated Content (UGC) in the real world. Furthermore, we present an end-to-end model designed to address the video repurposing challenge by jointly learning classification and regression tasks. Through extensive experiments and ablation studies, we demonstrate the effectiveness of the proposed model.